# Periodic classical trajectories and quantum scars in many-spin systems

### v2

@@ Abstract
To probe the limits of dynamic thermalisation, we numerically investigate the stability of exceptional periodic classical trajectories in rather generic chaotic many-body systems and explore a possible connection between these trajectories and exceptional nonthermal quantum eigenstates known as “quantum many-body scars”. The systems considered are chaotic spin chains with short-range interactions, both classical and quantum. On the classical side, the chosen periodic trajectories are such that all spins instantaneously point in the same direction, which evolves as a function of time. We find that the largest Lyapunov exponents characterising the stabillity of these trajectories have surprisingly strong and nontrivial dependencies on the interaction constants and chain lengths. In particular, we identify rather long spin chains, where the above periodic trajectories are Lyapunov-stable on many-body energy shells overwhelmingly dominated by chaotic motion. We show that the above phenomenology can be quantitatively described by connecting Lyapunov instabilities of translationally invariant periodic trajectories to irreducible representations of the translational symmetry group with well-defined wave vectors. We also find that instabilities around periodic trajectories in modestly large spin chains develop into a transient nearly quasiperiodic non-ergodic regime. In some cases, the lifetime of this regime is extremely long, which we interpret as a manifestation of Arnold diffusion in the vicinity of integrable dynamics. On the quantum side, we numerically investigate the dynamics of quantum states starting with all spins initially pointing in the same direction: these are the quantum counterparts of the initial conditions for the above periodic classical trajectories. Our investigation reveals the existence of quantum many-body scars for numerically accessible finite chains of spins 3/2 and higher. No evidence of quantum scars was observed for spin-1/2 chains, while spin-1 chains were found to be transitional in this respect. The dynamic thermalisation process dominated by quantum scars is shown to exhibit a slowdown in comparison with generic thermalisation at the same energy. Finally, we identify quantum signatures of the proximity to a classical separatrix of the periodic motion.
@@ I Introduction
The role and the properties of chaos in the process of dynamic thermalisation of many-body systems remain a subject of intense research that keeps resisting generic conjectures[1, 2, 3]. The open issues concern the transition from classical to quantum thermalisation[4, 5], as well as the thermalisation in purely classical systems as such[6, 7]. Central to those issues is the lack of a clear border between systems exhibiting and not exhibiting dynamic thermalisation to the Gibbs equilibrium. It is often the case that a seemingly exotic non-thermalising setting, in fact, sets a warning post for a much broader class of physical situations, where the thermalisation is absent or poor. In the latter case, the system may be thermalisable in principle but only on a rather long timescale[8, 9, 10, 11, 12, 13, 14, 15, 16]. The examples of nonthermalising/poorly thermalising behavior are: integrable systems [17, 18, 19, 20], glassy systems [21, 22], systems possibly exhibiting many-body localisation[23, 24, 25, 26, 27], periodically driven systems[28, 29, 30, 31, 32, 33, 34, 35, 36], and, finally, coming to the topic of this article, systems with quantum many-body scars [37, 38, 39].
Even when a system appears sufficiently “generic”, there is still a question[40, 41]: How does the character and the outcome of the thermalisation process depend on the initial conditions? One should keep in mind here that the process of dynamic thermalisation starts from a macroscopic nonequilibrium initial state of an isolated many-body system. Any macroscopic nonequilibrium state is extremely unlikely in comparison with the equilibrium one at the same energy. It is, therefore, intrinsically problematic to discard one nonequilibrium state as being more unlikely than another, as long as both states can be prepared experimentally.
@@
The present article is about thermalisation and chaos in relatively generic systems of interacting classical and quantum spins for somewhat exotic but experimentally realisable initial conditions where all spins initially point in the same direction. Classically, these initial conditions lead to short periodic trajectories. In what follows, we describe the fate of these periodic trajectories in two senses, namely: fate following small deviations from these trajectories in purely classical systems, and also fate in the sense of transitioning from classical to quantum dynamics. Our original motivation was to explore whether periodic trajectories in the classical limit entail a quantum-many-body-scar behavior, and, indeed we report here a positive numerical evidence of quantum scars for quantum spins $S\geq 3/2$. However, the investigation also revealed a number of surprising properties on the purely classical side, such as the non-trivial dependence of the Lyapunov stability of the periodic orbits on the chain length, including fairly long spin chains with Lyapunov-stable periodic orbits on otherwise strongly chaotic energy shells. We also discovered that unstable periodic motion in finite spin systems evolves to exhibit a transient nearly quasiperiodic non-ergodic regime. In several cases, this regime has an anomalously long lifetime, which we interpret as a manifestation of Arnold diffusion[42]. Various classical regimes investigated in this work are schematically summarised in Fig. 1.
@@ II.1 General aspects
The practical relevance of the KAM theorem to the thermalisation in many-body systems is sometimes viewed skeptically based on the expectation that the parameter range where the KAM tori survive decreases exponentially with the number of degrees of freedom in the system, and, even when some of them do survive, they do not suppress ergodicity of a typical phase space trajectory. Yet, contrary to the latter view, the findings of the present work indicate the surprising relevance of the KAM-tori even to the dynamics of strongly nonintegrable systems. These findings include the existence of Lyapunov-stable periodic trajectories (Sections IV) and the existence of a nearly quasiperiodic regime (Sections V) – both occurring in finite but large systems on energy shells dominated by strong chaotic motion.
@@
When it comes to the transition from classical to quantum thermalisation, one of the primary challenges here is the fact that the classical description is based on the notions of ergodicity and chaos of phase space trajectories, while quantum-mechanically the very notion of a phase trajectory is not definable – consequence of the Heisenberg uncertainty principle not allowing one to simultaneously determine the coordinates and momenta of particles with infinite precision. On the quantum side, chaotic instabilities and ergodicity become detached from each other. The common indicators of “quantum chaos”, such as the statistics of energy level spacings [51, 52] or eigenstate thermalisation [53, 54, 55, 56, 57, 58], involve the global properties of many-body energy eigenstates, and therefore, characterise their ergodicity. At the same time, recent research on out-of-time-ordered correlators (OTOCs) indicates[59, 60, 61, 62, 63] that chaotic instabilities in response to small deviations of initial conditions are, strictly speaking, absent in quantum systems with bounded local operators (e.g. spins 1/2) and finite-range interactions even in the thermodynamic limit. (Here, we sidestep the models with infinite-range interactions, where chaos and thermalisation properties may be rather special[64, 65, 66, 67].) Systems with finite-range interactions, however, may still exhibit a transient exponential OTOC growth representative of the classical chaotic limit, in particular, for lattices of spins $S>1/2$ [60, 62]. A similar condition, namely, $S\geq 3/2$, to observe the effect of classical periodic trajectories on quantum dynamics is reported in Section VI of the present work. The readers are referred to Refs.[68, 69, 70, 71, 72, 73, 74] for various complementary aspects of the correspondence between classical and quantum spin dynamics.
@@
Other efforts to identify the indicators and the mechanisms of the proper quantum thermalisation targeted the properties of the Hamiltonian matrix elements[75, 76], the role of quantum typicality[77, 78, 79, 80, 81, 82], generic features of the relaxation behavior[83, 84, 85, 86, 87], the frequency spectra of the observables[88, 89], adiabatic eigenstate deformations[90, 91], quantum scrambling[92, 93, 94, 95, 96, 97, 98] and, more recently, the universal operator growth[89, 99, 100, 101]. All the above properties are subject to caution due to possible generic existence of exceptional quantum-scar states associated with the periodic classical orbits of the kind considered in this work.
@@
Another frontier of the thermalisation research is the behavior of large but finite systems. The importance of this frontier is that macroscopic systems that exhibit nontrivial thermalisation behavior are often those that also exhibit fragmentation into weakly coupled nanoscale subsystems – e.g. large molecules[102] or clusters of different phases[103]. Each of such subsystems can in turn exhibit an unusual dynamics associated with its finite size[29, 33, 104, 34]. The stationary state of a finite system can realistically deviate from the Gibbs equilibrium, because the number of degrees of freedom and/or the number of quantum levels is insufficiently large. For finite classical systems, the concerns related to the KAM theorem also become more plausible. The present work contributes to the above agenda by showing that the stability of periodic trajectories in finite classical spin chains exhibits a strong non-trivial dependence on the chain length including strongly chaotic and non-chaotic cases, and that unstable periodic motion in finite systems decays into a transient nearly quasiperiodic regime.
@@ II.2 Quantum scars
A recent addition to the zoo of poorly thermalising settings is the case of the so-called “quantum many-body scars” associated with anomalously slow relaxation of quantum lattice systems that was first observed experimentally in Ref.[38] and then interpreted theoretically in Ref.[39] in terms of the “PXP” quantum model for a spin-1/2 chain. It was found that the PXP model possesses atypical eigenstates inconsistent with eigenstate thermalisation[53, 54], and, at the same time, these atypical eigenstates are prominently present in the initial nonequilibrium state of the experimentally observed system of Ref.[38]. Eigenstate thermalisation hypothesis (ETH) stipulates that the quantum expectation values of a physical observable in a normally thermalising system remain almost the same for all energy eigenstates within a given energy window. The atypical eigenstates of the PXP model named “quantum scars” stood out in this respect. A similar kind of ETH violation was, actually, reported earlier in Ref. [37] without reference to quantum scars. The above findings were followed a large body of research [105, 106, 107, 108, 109, 110, 111, 112, 113].
@@
Let us recall here that the concept of quantum scars as such was initially introduced by Heller[114] in the eighties for one particle moving in a chaotic two-dimensional billiard: it referred to atypical quantum eigenstates corresponding to classical periodic trajectories with short periods. When a classical periodic trajectory is sufficiently short, then the quasiclassical limit of the quantum dynamics implies the onset of quantum interference corrections associated with the trajectory closing onto itself before the chaotic Lyapunov instability suppresses such interference, and this corrections, in turn, amplify the probability of the corresponding eigenstate to remain localized around the classical trajectory as opposed to being evenly spread over the entire system which the typical eigenstates do. Numerous examples of low-dimensional quantum scars have been investigated since then – see, e.g., [115, 116]. One missing link between the original low-dimensional quantum scars and the recently discovered quantum many-body scars was that the latter were not connected to periodic classical trajectories. In other words, the quantum many-body scars as discovered were rather “distant relatives” to the quantum scars known from the physics of billiards. This issue of missing periodic trajectories was addressed in [117, 118], although not in the classical limit but rather in the parameter space of quantum matrix product states.
When completing this article, we discovered Ref. [113] that was pursuing a very similar agenda but in the context of the Bose-Hubbard model — looking for quantum scars around a periodic orbit of a classical mean-field solution. Ref. [113] and the present work appear to largely agree on their conclusions, while arriving to them by rather complementary routes. We further note in this regard that most of the findings of the present paper were first reported in the Ph.D. thesis [119].
@@ II.3 Outline of the article
In Section VI, we numerically investigate chains of quantum spins $1/2$, 1, $3/2$ and 2 with varying chain lengths looking for the signatures of quantum scars corresponding to the above classical periodic orbits. We identify such signatures for spins $3/2$ and 2. The signatures include the slowdown of spin relaxation and the violation of the eigenstate thermalisation hypothesis. We also report on the quantum signatures of the proximity to the classical separatrix between librations and rotations.
@@ III Hamiltonian and initial conditions
The particular form of the Hamiltonian III is motivated by its robust non-integrability in the quantum case as evidenced by the absence of the significant finite-size effects in the energy level spacing statistics (see Fig. 15 below).
We investigate one special type of nonequilibrium initial conditions corresponding to all spins polarised in the same direction. Specifically, we choose them polarised along the $z$-direction, which, according to Hamiltonian (III), implies that the initial energy (energy expectation value in the quantum case) is equal to zero, which, in turn, corresponds to the infinite temperature. The latter implies that the initial state belongs to the energy shell of maximum entropy (maximum density of states quantum mechanically); such a choice minimizes the influence of the finite-size effects on our numerical results.

### v3

@@ Abstract
To probe the limits of dynamic thermalization, we numerically investigate the stability of exceptional periodic classical trajectories in chaotic many-spin systems and explore a possible connection between these trajectories and exceptional nonthermal quantum eigenstates known as “quantum many-body scars”. The systems considered are chaotic spin chains with short-range interactions, both classical and quantum. On the classical side, the chosen periodic trajectories are such that all spins instantaneously point in the same direction, which evolves as a function of time. We find that the largest Lyapunov exponents characterising the stabillity of these trajectories have surprisingly strong and nontrivial dependencies on the interaction constants and chain lengths. In particular, we identify rather long spin chains, where the above periodic trajectories are Lyapunov-stable on many-body energy shells overwhelmingly dominated by chaotic motion. We show that the above phenomenology can be quantitatively described by connecting Lyapunov instabilities of translationally invariant periodic trajectories to irreducible representations of the translational symmetry group with well-defined wave vectors. We also find that instabilities around periodic trajectories in modestly large spin chains develop into a transient nearly quasiperiodic non-ergodic regime. In some cases, the lifetime of this regime is extremely long, which we interpret as a manifestation of Arnold diffusion in the vicinity of integrable dynamics. On the quantum side, we numerically investigate the dynamics of quantum states starting with all spins initially pointing in the same direction: these are the quantum counterparts of the initial conditions for the above periodic classical trajectories. Our investigation reveals the existence of quantum many-body scars for numerically accessible finite chains of spins 3/2 and higher. No evidence of quantum scars was observed for spin-1/2 chains, while spin-1 chains were found to be transitional in this respect. The dynamic thermalization process dominated by quantum scars is shown to exhibit a slowdown in comparison with generic thermalization at the same energy. Finally, we identify quantum signatures of the proximity to a classical separatrix of the periodic motion.
@@ I Introduction
The role and the properties of chaos in the process of dynamic thermalization of many-body systems remain a subject of intense research that keeps resisting generic conjectures[1, 2, 3]. The open issues concern the transition from classical to quantum thermalization[4, 5], as well as the thermalization in purely classical systems as such[6, 7]. Central to those issues is the lack of a clear border between systems exhibiting and not exhibiting dynamic thermalization to the Gibbs equilibrium. It is often the case that a seemingly exotic non-thermalising setting, in fact, sets a warning post for a much broader class of physical situations, where the thermalization is absent or poor. In the latter case, the system may be thermalisable in principle but only on a rather long timescale[8, 9, 10, 11, 12, 13, 14, 15, 16]. The examples of nonthermalising/poorly thermalising behavior are: integrable systems [17, 18, 19, 20], glassy systems [21, 22], systems possibly exhibiting many-body localisation[23, 24, 25, 26, 27], periodically driven systems[28, 29, 30, 31, 32, 33, 34, 35, 36], and, finally, coming to the topic of this article, systems with quantum many-body scars [37, 38, 39].
Even when a system appears sufficiently “generic”, there is still a question[40, 41]: How does the character and the outcome of the thermalization process depend on the initial conditions? One should keep in mind here that the process of dynamic thermalization starts from a macroscopic nonequilibrium initial state of an isolated many-body system. Any macroscopic nonequilibrium state is extremely unlikely in comparison with the equilibrium state at the same energy. It is, therefore, intrinsically problematic to discard one nonequilibrium state as being more unlikely than another, as long as both states can be prepared experimentally.
@@
The present article is about thermalization and chaos in relatively generic systems of interacting classical and quantum spins for somewhat exotic but experimentally realisable initial conditions where all spins initially point in the same direction. Classically, these initial conditions lead to short periodic trajectories. In what follows, we describe the fate of these periodic trajectories in two senses, namely: fate following small deviations from these trajectories in purely classical systems, and also fate in the sense of transitioning from classical to quantum dynamics. Our original motivation was to explore whether periodic trajectories in the classical limit entail a quantum-many-body-scar behavior, and, indeed we report here a positive numerical evidence of quantum scars for quantum spins $S\geq 3/2$. However, the investigation also revealed a number of surprising properties on the purely classical side, such as the non-trivial dependence of the Lyapunov stability of the periodic orbits on the chain length, including fairly long spin chains with Lyapunov-stable periodic orbits on otherwise strongly chaotic energy shells. We also discovered that unstable periodic motion in finite spin systems evolves to exhibit a transient nearly quasiperiodic non-ergodic regime. In several cases, this regime has an anomalously long lifetime, which we interpret as a manifestation of Arnold diffusion[42]. Various classical regimes investigated in this work are schematically summarised in Fig. 1.
@@ II.1 General aspects
The practical relevance of the KAM theorem to the thermalization in many-body systems is sometimes viewed skeptically based on the expectation that the parameter range where the KAM tori survive decreases exponentially with the number of degrees of freedom in the system, and, even when some of them do survive, they do not suppress ergodicity of a typical phase space trajectory. Yet, contrary to the latter view, the findings of the present work indicate the surprising relevance of the KAM-tori even to the dynamics of strongly nonintegrable systems. These findings include the existence of Lyapunov-stable periodic trajectories (Sections IV) and the existence of a nearly quasiperiodic regime (Sections V) – both occurring in finite but large systems on energy shells dominated by strong chaotic motion.
@@
When it comes to the transition from classical to quantum thermalization, one of the primary challenges here is the fact that the classical description is based on the notions of ergodicity and chaos of phase space trajectories, while quantum-mechanically the very notion of a phase trajectory is not definable – consequence of the Heisenberg uncertainty principle not allowing one to simultaneously determine the coordinates and momenta of particles with infinite precision. On the quantum side, chaotic instabilities and ergodicity become detached from each other. The common indicators of “quantum chaos”, such as the statistics of energy level spacings [51, 52] or eigenstate thermalization [53, 54, 55, 56, 57, 58], involve the global properties of many-body energy eigenstates, and therefore, characterise their ergodicity. At the same time, recent research on out-of-time-ordered correlators (OTOCs) indicates[59, 60, 61, 62, 63] that chaotic instabilities in response to small deviations of initial conditions are, strictly speaking, absent in quantum systems with bounded local operators (e.g. spins 1/2) and finite-range interactions even in the thermodynamic limit. (Here, we sidestep the models with infinite-range interactions, where chaos and thermalization properties may be rather special[64, 65, 66, 67].) Systems with finite-range interactions, however, may still exhibit a transient exponential OTOC growth representative of the classical chaotic limit, in particular, for lattices of spins $S>1/2$ [60, 62]. A similar condition, namely, $S\geq 3/2$, to observe the effect of classical periodic trajectories on quantum dynamics is reported in Section VI of the present work. The readers are referred to Refs.[68, 69, 70, 71, 72, 73, 74] for various complementary aspects of the correspondence between classical and quantum spin dynamics.
@@
Other efforts to identify the indicators and the mechanisms of the proper quantum thermalization targeted the properties of the Hamiltonian matrix elements[75, 76], the role of quantum typicality[77, 78, 79, 80, 81, 82], generic features of the relaxation behavior[83, 84, 85, 86, 87], the frequency spectra of the observables[88, 89], adiabatic eigenstate deformations[90, 91], quantum scrambling[92, 93, 94, 95, 96, 97, 98] and, more recently, the universal operator growth[89, 99, 100, 101]. All the above properties are subject to caution due to possible generic existence of exceptional quantum-scar states associated with the periodic classical orbits of the kind considered in this work.
@@
Another frontier of the thermalization research is the behavior of large but finite systems. The importance of this frontier is that macroscopic systems that exhibit nontrivial thermalization behavior are often those that also exhibit fragmentation into weakly coupled nanoscale subsystems – e.g. large molecules[102] or clusters of different phases[103, 104]. Each of such subsystems can in turn exhibit an unusual dynamics associated with its finite size[29, 33, 105, 34]. The stationary state of a finite system can realistically deviate from the Gibbs equilibrium, because the number of degrees of freedom and/or the number of quantum levels is insufficiently large. For finite classical systems, the concerns related to the KAM theorem also become more plausible. The present work contributes to the above agenda by showing that the stability of periodic trajectories in finite classical spin chains exhibits a strong non-trivial dependence on the chain length including strongly chaotic and non-chaotic cases, and that unstable periodic motion in finite systems decays into a transient nearly quasiperiodic regime.
@@ II.2 Quantum scars
A recent addition to the zoo of poorly thermalising settings is the case of the so-called “quantum many-body scars” associated with anomalously slow relaxation of quantum lattice systems that was first observed experimentally in Ref.[38] and then interpreted theoretically in Ref.[39] in terms of the “PXP” quantum model for a spin-1/2 chain. It was found that the PXP model possesses atypical eigenstates inconsistent with eigenstate thermalization[53, 54], and, at the same time, these atypical eigenstates are prominently present in the initial nonequilibrium state of the experimentally observed system of Ref.[38]. Eigenstate thermalization hypothesis (ETH) stipulates that the quantum expectation values of a physical observable in a normally thermalising system remain almost the same for all energy eigenstates within a given energy window. The atypical eigenstates of the PXP model named “quantum scars” stood out in this respect. A similar kind of ETH violation was, actually, reported earlier in Ref. [37] without reference to quantum scars. The above findings were followed a large body of research [106, 107, 108, 109, 110, 111, 112, 113, 114].
@@
Let us recall here that the concept of quantum scars as such was initially introduced by Heller[115] in the eighties for one particle moving in a chaotic two-dimensional billiard: it referred to atypical quantum eigenstates corresponding to classical periodic trajectories with short periods. When a classical periodic trajectory is sufficiently short, then the quasiclassical limit of the quantum dynamics implies the onset of quantum interference corrections associated with the trajectory closing onto itself before the chaotic Lyapunov instability suppresses such interference, and these corrections, in turn, amplify the probability of the corresponding eigenstate to remain localized around the classical trajectory as opposed to being evenly spread over the entire system which the typical eigenstates do. Numerous examples of low-dimensional quantum scars have been investigated since then – see, e.g., [116, 117]. Yet, a missing link between the original low-dimensional quantum scars and the recently discovered quantum many-body scars was that the latter were not connected to periodic classical trajectories. In other words, the quantum many-body scars as discovered were rather “distant relatives” to the quantum scars known from the physics of billiards. This issue of missing periodic trajectories was addressed in [118, 119], although not in the classical limit but rather in the parameter space of quantum matrix product states.
When completing this article, we discovered that a very similar agenda was pursued for two-spin spin systems in Refs. [120, 121], for a two-site Bose-Hubbard model in Ref. [122] and for a many-site Bose-Hubbard model in Ref. [114]. The above references and the present work appear to largely agree on their conclusions, while arriving to them by rather complementary routes. We further note in this regard that most of the findings of the present paper were first reported in the Ph.D. thesis [123].
@@ II.3 Outline of the article
In Section VI, we numerically investigate chains of quantum spins $1/2$, 1, $3/2$ and 2 with varying chain lengths looking for the signatures of quantum scars corresponding to the above classical periodic orbits. We identify such signatures for spins $3/2$ and 2. The signatures include the slowdown of spin relaxation and the violation of the eigenstate thermalization hypothesis. We also report on the quantum signatures of the proximity to the classical separatrix between librations and rotations.
@@ III Hamiltonian and initial conditions
The particular form of the Hamiltonian III is motivated by its robust non-integrability in the quantum case. In particular, choosing $h$ on the order of $J$ ensures that the chain with spins 1/2 is far from the integrable limits of either $h=0$ or $J=0$. In addition, the confinement physics often obscuring thermalization in finite systems [9, 124, 125] has been avoided by introducing both interaction terms $S_{i}^{x}S_{i+1}^{x}$ and $S_{i}^{y}S_{i+1}^{y}$. Finally, we have avoided symmetry with respect to the exchange $\{S_{i}^{x}\leftrightarrow S_{i}^{y}\}$ by having the $yy$-coupling twice that of $xx$. (See Section VI for further discussion.)
We investigate one special type of nonequilibrium initial conditions corresponding to all spins polarised in the same direction. Specifically, we choose them polarised along the $z$-direction, which, according to Hamiltonian (III), implies that the initial energy $E$ (energy expectation value in the quantum case) is equal to zero, which, in turn, corresponds to the infinite temperature in the sense of canonical ensemble. Indeed, since the infinite-temperature classical probability distribution is uniform in the classical phase space, while the infinite-temperature quantum density matrix is proportional to the unit matrix in the quantum Hilbert space, the average energy in the both cases can be written as
$E_{\infty}=\langle\mathcal{H}\rangle=-\sum\limits^{L}_{i=1}$ $\left(J\langle S_{i}^{x}S_{i+1}^{x}\rangle+2J\langle S_{i}^{y}S_{i+1}^{y}\rangle\right)$
$+\sum\limits^{L}_{i=1}\left(h\langle S_{i}^{x}\rangle+h\langle S_{i}^{y}\rangle\right)=0.$ (2)
Here delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over all possible values of spin projections, quantum or classical, which implies that $\langle S_{i}^{\alpha}\rangle=0$ and $\langle S_{i}^{\alpha}S_{i+1}^{\alpha}\rangle=0$. Further noting the equivalence between the canonical and the microcanonical ensembles in the limit $L\to\infty$, we conclude that an initial spin state with zero energy belongs to the energy shell of maximum entropy (maximum density of states quantum mechanically).
@@
We note that all couplings in the Hamiltonian (III) were chosen to involve only $x$ and $y$ spin projections to ensure that the initial state fully polarized along the $z$-axis has zero energy and hence lies in the middle of the energy spectrum, which is important for reducing the finite-size effects on spin dynamics.